equations by elimination? -4x-2y=-12
4x+8y=-24

Found 2 solutions by Alan3354, sofiyac:
Answer by Alan3354(69443)

Add the 2 equations.

Answer by sofiyac(983)

add the two equations and your x terms add to zero, so you get
6y=-36 divide each side by 6 and you get
y=-6 now plug that into either of the equations and solve for x
4x+8*-6=-24
4x-48=-24
4x=24
x=6
(6,-6)